OpenStudy (anonymous):

how many different committees can be formed from 10 teachers and 41 students if the committee consists of 3 teachers and 2 students

OpenStudy (loser66):

\(\left(\begin{matrix}10\\3\end{matrix}\right)\)\(\left(\begin{matrix}41\\2\end{matrix}\right)\)